OpenStudy (princessaurora):

*Trigonometric Functions* What is the solution of this equation over the interval [0,2pi]: cos(x)tan(x)-1/2=0? *help is appreciated and will give medal*

Nnesha (nnesha):

what's the reciprocal of tan ?

OpenStudy (anonymous):

good job princess

OpenStudy (princessaurora):

sec?

Nnesha (nnesha):

no cos =1/sec and sin=1/csc tan =what ? what's the definition of tan how would you write tan in terms of sin cos ?

OpenStudy (anonymous):

what ? sec ? lol

OpenStudy (anonymous):

Cas1 is a skank retricewhore

OpenStudy (anonymous):

lol no one knows

OpenStudy (anonymous):

lol be quiet @chafrees1 i love u bruh

OpenStudy (princessaurora):

tan x° = Opposite / Adjacent?

Nnesha (nnesha):

relation between trig functions 

OpenStudy (princessaurora):

Thank you I see now.

Nnesha (nnesha):

tan = sin/cos \[\rm \cos(x)\color{ReD}{ \tan(x)}-\frac{1}{2}=0\] replace tan with sin/cos and then simplify it

OpenStudy (anonymous):

your welcome

OpenStudy (princessaurora):

cos(x)sin(x)/cos(x)-1/2=0?

Nnesha (nnesha):

\[\huge\rm \cos(x)*\frac{\sin(x)}{\cos(x)}-\frac{1}{2}=0\] which is same as \[\frac{ \cos(x)*\sin(x) }{ \cos(x)}-\frac{1}{2}=0 \]

OpenStudy (princessaurora):

Ok I see

Nnesha (nnesha):

simplify that

OpenStudy (princessaurora):

x=π6±2πn,5π6±2πn?

Nnesha (nnesha):

hamm what's that ? what would be the next step ? \[\frac{ \cos(x)*\sin(x) }{ \cos(x)}-\frac{1}{2}=0 \]

OpenStudy (princessaurora):

to multiply cos(x) and sin(x)?

Nnesha (nnesha):

i already did part now cos(x) times sin (x) over cos(x) = ??

OpenStudy (princessaurora):

I don't know I'm stuck on this part

Nnesha (nnesha):

let cos(x) = y and sin(x) = z \[\color{Red}{\frac{ \cos(x)*\sin(x) }{ \cos(x)}}-\frac{1}{2}=0 \] \[\color{Red}{\frac{ y*z}{ y}}-\frac{1}{2}=0 \]

Nnesha (nnesha):

\[\rm \frac{ y*z }{ y }=??\]

OpenStudy (princessaurora):

yz/y?

Nnesha (nnesha):

ye correct what else ? simplify that

OpenStudy (princessaurora):

z is all that's left?

Nnesha (nnesha):

right and we supposed that z=sin(x) same thing let cos(x) = y and sin(x) = z \[\color{Red}{\frac{ \cancel{\cos(x)}*\sin(x) }{ \cancel{\cos(x)}}}-\frac{1}{2}=0 \] \[\sin(x)-\frac{1}{2}=0\] now jsut solve for sin(x)

OpenStudy (princessaurora):

sin(x) is z?

OpenStudy (princessaurora):

so z=1/2?

Nnesha (nnesha):

well i just replaced sin(x) with z to make it looks easy

Nnesha (nnesha):

yes right so sin(x) = 1/2 now u can use unit circle to find exact solution from 0 to 2pi

Nnesha (nnesha):

 (x,y) order pair where x-coordinate =cos and y-coordinate = sin

OpenStudy (princessaurora):

(1,0)?

Nnesha (nnesha):

no sin(x)=1/2 you have to find x value at what angle sin(x) is 1/2 ?